Write the point slope form of the equation of the line that passes through the points (-8,2) and (1,-4)
Found 2 solutions by math_helper, Ramzan:
Answer by math_helper(2461)

Point slope form is:
y-y1 = m(x-x1)

m = slope = change_in_y/change_in_x = (-4-2)/(1-(-8) = -6/9 = -2/3

Equation of line in point slope form is: y-2 = -(2/3)(x + 8)
--
You can use either point.
If you do use the other point, you should be able to rearrange the resulting equation to be exactly the same as the one above.

Answer by Ramzan(1)

Point slope form is:
y-y1 = m(x-x1)

m = slope = change_in_y/change_in_x = (-4-2)/(1-(-8) = -6/9 = -2/3

Equation of line in point slope form is: y-2 = -(2/3)(x + 8)